OpenStudy (anonymous):

h^2-6h=72 how do you find h?

OpenStudy (johnweldon1993):

We can either factor, or complete the square

OpenStudy (anonymous):

the answer should be 12 but I need to work it out and I don't know how to

OpenStudy (anonymous):

First set it equal to zero then you can use @johnweldon1993 method or even the quadratic formula

OpenStudy (anonymous):

can you add the h^2 and the -6h

OpenStudy (johnweldon1993):

To factor, subtract 72 from both sides \[\large h^2 - 6h - 72\] Now find the factors of -72 that also add up to -6 ***also, there will be 2 answers***

OpenStudy (johnweldon1993):

So lets see 1 and 72 2 and 36 3 and 24 4 and 18 6 and 12 <--- oh look We know that 6 and 12 can somehow be combined to add to -6 \[\large (x - 12)(x + 6)\] This shows us that both 12 and -6 can be answers to our equation

OpenStudy (anonymous):

well there is suppose to only be one answer so how do I work it out so I can only get one asnwer

OpenStudy (johnweldon1993):

You cannot, there will always be 2 answers to a quadratic equation, perhaps the teacher may only want 1 answer recorded, however this is technically wrong since there WILL be 2 results

OpenStudy (johnweldon1993):

We can use another method to show that too, completing the square \[\large h^2 - 6h = 72\] Take half the coefficient of the 'h' (which is -6) divide it by 2 (so -3) and then square that result (so that will come out to 9) Now add that to both sides \[\large h^2 - 6h + 9 = 72 + 9\] \[\large h^2 - 6h + 9 = 81\] Turn the left hand side into a perfect square \[\large (h - 3)^2 = 81\] take the square root of both sides \[\large h - 3 = \pm 9\] This will lead to 2 equations \(\large h - 3 = 9\) and \(\large h - 3 = -9\) solving both f those we get \(\large h = 12\) and \(\large h = -6\)

OpenStudy (anonymous):

the area of a triangle is given by the equation h^2-6h=72 where h is the height of the triangle. what is the value of h? there should only be one answer

OpenStudy (johnweldon1993):

So the teacher is justified in asking for only 1 answer, we cannot have a negative area, the item would be in some parallel universe for that to happen, so 12 would be the answer recorded there

OpenStudy (anonymous):

ok so if 12 is the answer how did you get 12